a stone is thrown vertically upwards with an initial velocity of 30 m/s. The time taken for the stone to rise to its maximum height is

Grade:11

2 Answers

Yashwant yadav
6 years ago
Basically ,by v=u-gt (-g represents opposite acceleration of object with g) ..now V=0 (at max) ,u=30,g =10 then....we get.t=3sec
kumar kanishk
6 years ago
given
u=30m/s
we know that at max height velocity becomes zero.
so from v = u + at
             0 = 30 – 10t(a=-g =10)
            ie 30=10t
       so t=3s